# Contrast Sets for Evaluating Language-Guided Robot Policies

### v1

@@ 6.1 Experiment Design
We design a pseudo studio apartment environment which is populated with furniture similar to those found in simulation. To ensure ecological validity of test instances, we recruited five participants to design five furniture setups, each of which they annotated with a navigation instruction similar to those in Figure 4. More details on this protocol and the training of the model can be found in the Appendix. Each test instance $x$ and outcome $y$ is defined by the scene $s$ based on the furniture arrangement and the robot’s pose, and the expected behavior $b$ of the robot, described by the language instruction $l$ consisting of two subgoals to be reached.
Perturbations. Figure 4 showcases the defined perturbation functions. We define language-only perturbations $\Delta L$ to simply change the wording of the instruction but preserving the meaning. $\Delta LB$ changes the final goal in the text instruction, so the first half of the robot’s behavior to the first goal is the same, but the latter half changes. $\Delta S$ moves an object around in the scene such that the expected trajectory of the robot is still the same. For example, passive objects the robot is not meant to interact with may change positions as long as they do not change the interpretation of either language subgoal. Lastly, $\Delta SB$ moves an object around in the scene such that the expected trajectory of the robot is different. This perturbation either moves the goal object such that the trajectory must be different, or an object is moved in front of the robot to impede its originally-intended trajectory.
Metrics. We define the cost metric $C$ as the distance objects were moved in the scene. Though the objects moved can vary from small tables to large couches, we do not use the weight of the item moved in our cost. Since there are two subgoals, the metric $M$ is the average progress towards goal. which is 50% for reaching the first goal of 100% for reaching both goals. Unlike simulation, real robot test instances cannot be exactly replicated, so we evaluate each test instance three times. With four perturbation strategies executed on each of the five original test instances, we have a total of 20 new test instances that are executed three times, for strategy size $|\Gamma({X})|=75$.
@@ 6.2 Results and Discussion
Modifying the scene and robot malfunctions—such as wheels breaking, hardware needing to be maintained, and software glitches—all add to the time and costs it take to evaluate robot policies. Experimenters have limited time, which motivates the need for more efficient evaluation.
Contrast set perturbations more enable experiments for less work. We found that the cost of the standard evaluation set is 8.8 meters. Since the contrast set evaluation is a superset of the standard test set, the cost of the additional perturbations adds around 1 meter of furniture moved. This additional work of 1 meter of scene perturbations allows the experimenter to execute 20 more trials.
@@
Contrast sets can estimate the full test set with less cost. Figure 5 demonstrates that contrast set evaluation converges to a similar success rate as standard evaluation while requiring less total effort. This result indicates that using contrast set evaluation allows experimenters to measure the robot policy’s performance with significantly less effort.
Contrast sets can provide insights into a policy’s linguistic and visual understanding. Similar to results in the simulation results, our specific perturbations can show a policy’s sensitivity to specific axes. We find that standard evaluation has a 43% average SR, but $\Delta LB$ and $\Delta VB$ subsets have 13% higher average success rates than without expected behavior changes. For $\Delta LB$, the final goal of the test instance was changed. Though the first half of the robot’s trajectory remained the same, this robot’s improvement on this perturbation may help experimenters quickly discover which objects and locations a policy struggles with. Similarly, $\Delta VB$ moves a single object that is referred to in the instruction without modifying the scene. This change can help characterize which objects the policy may struggle identifying. Though our perturbations for this experiment are relatively general they can already provide experimenters with intuitions on policy performance. More specific perturbations can highlight regions of the test domain in which the policy is particularly effective and where it may need further improvement.
@@ 7 Conclusion and Limitations
A language-guided robot can be tasked with potentially thousands of tasks across arbitrary environments. To evaluate such a robot, experimenters ideally construct i.i.d. test sets that can probe a robot’s performance across the domain of possible test instances. Though these i.i.d test instances likely cover the domain of test instances effectively, they may require the experimenter to shuffle the room around extensively. This process requires more experimenter effort. In this work, we propose contrast sets as an approach for evaluating language-guided robot policies. We find that evaluating contrast sets around test instances provides insights into policy robustness and sensitivity with respect to input language and visual scene without compromising an estimate of the full test set performance and with less experimenter cost. We designed a simulated manipulation experiment and a real-world vision-and-language navigation experiment, and found that contrast sets allow experimenters to run more evaluations with less effort. We argue that contrast set evaluations offer higher fidelity than small-scale, real robot demonstrations while not requiring the industry-level resources for large-scale, deployed evaluations.
While we focused primarily on environments where language instructions specify greater details on how the robot should interact with the environment, we intend to investigate larger manipulation-focused datasets such as DROID [34] and OXE [35]. As these datasets become larger and models become more expressive, efficiently evaluating the thousands of tasks these datasets enable will become important. Our work assumes the existence of a test set and that it is representative of the test manifold. Unlike other works, we are not trying to estimate performance outside of the manifold, and we assume that contrast set perturbations would generate variants that are in the domain of the problem. Typically, these test sets are not well-defined in advance and future work can investigate which types of test instance perturbations can allow experimenters to efficiently evaluate the capabilities of more open-ended robot systems.
@@ Acknowledgments
The authors would like to thank Elle Szabo for building the early infrastructure for the physical robot experiments.
Appendix A Language-Table Simulator Experiments
@@ A.3 Additional Results
Appendix B VLN-CE Evaluation on a Physical Robot

### v2

@@ 6.1 Experiment Design
We design a pseudo studio apartment environment populated with furniture similar to categories in simulation. To ensure ecological validity of test instances, we recruited five participants to design five furniture setups, each of which they annotated with a navigation instruction as shown in Figure 4. Each test instance $x$ and outcome $y$ is defined by the scene $s$ based on the furniture arrangement and the robot’s pose, and the expected behavior $b$ of the robot, described by the language instruction $l$ consisting of two subgoals to be reached. More details on this protocol for ensuring ecological validity and model training can be found in Appendix C.
Perturbations. We define four perturbation functions as shown in Figure 4. We define language-only perturbations $\Delta L$ to simply change the wording of the instruction but preserving the meaning. If a language instruction tells the policy to go to a bed then to the couch, then this perturbation simply rephrases the instruction. $\Delta LB$ changes the final goal in the text instruction, so the first half of the robot’s behavior to the first goal is the same, but the latter half changes. $\Delta S$ moves an object around in the scene such that the expected trajectory of the robot is still the same. For example, passive objects the robot is not meant to interact with may change positions as long as they do not change the interpretation of either language subgoal. Lastly, $\Delta SB$ moves an object around in the scene such that the expected trajectory of the robot is different. This perturbation either moves the goal object such that the trajectory must be different, or an object is moved in front of the robot to impede its originally-intended trajectory.
Metrics. We define the cost metric $C$ as the distance objects were moved in the scene. The metric $M$ is the average progress toward two subgoals: 50% for reaching the first and 100% for reaching both. We evaluate each test instance three times. Four perturbation strategies applied to five original test instances yield 20 new instances Executing these 20 test instances and the original 5 test instances three times results in a strategy set size of $|\Gamma({X})|=75$.
@@ 6.2 Results and Discussion
Modifying the scene, robot malfunctions, and maintenance all add to the cost of evaluating robots. We find that contrast sets reduce the effort required for evaluating robots in the real world.
Contrast set perturbations enable more experiments for less work. The contrast set consists of 4 perturbation functions. Based on Figure 8 in the Appendix,the average cost per instance in the standard evaluation set is 8.8 meters, while generating 4 perturbed trials from a single i.i.d. test instance costs about 1 meter. With 5 i.i.d. test instances from the standard evaluation set, the additional cost of 1 meter from perturbations allows one to execute $4*5=20$ more trials.
Contrast sets can estimate the full test set with less cost. Figure 5 demonstrates that contrast set evaluation converges to a similar success rate as standard evaluation while requiring less effort, indicating that this evaluation allows experimenters to efficiently measure a policy’s performance.
@@
Contrast sets can provide insights into a policy’s linguistic and visual understanding. Similar to results in the simulation results, our specific perturbations can show a policy’s sensitivity to specific axes. We find that $\Delta L$ and $\Delta S$ is lower in performance than that of the full contrast set, especially compared to the higher performing contrast sets. For example, we found that instruction “Go pass by the lamp and the plants, keep going and stop when you reach the bookshelf” performed with an average progress to goal of 100% over three runs, while the reworded version in the same setting achieved 33%. We found this trend to continue, and it directly highlights potential issues in the language encoder, as the language representations are likely not well aligned. Though our perturbations for this experiment are relatively general, more specific perturbations can highlight regions of the test domain where a policy is particularly effective and where it may need improvement.
@@ 7 Conclusion and Limitations
A language-guided robot can be tasked with potentially thousands of tasks across arbitrary environments. To evaluate such a robot, experimenters ideally construct i.i.d. test sets that effectively cover the domain of possible test instances. However, creating these sets often require extensive rearrangement of the environment, increasing experimenter effort. In this work, we propose contrast sets as an approach for evaluating language-guided robot policies. We find that evaluating these contrast sets provides insights into policy robustness and sensitivity. Additionally, contrast sets are able to estimate the full evaluation set performance while maintaining low experimenter cost. We conducted simulated manipulation and real-world vision-and-language navigation experiments and found that contrast sets enable experimenters to run more evaluations with less effort. We argue that contrast set evaluations offer higher fidelity than small-scale, real robot demonstrations while not requiring the industry-level resources for large-scale, deployed evaluations.
While we focused primarily on environments where language instructions specify greater details on how the robot should interact with the environment, we intend to investigate larger manipulation-focused datasets such as DROID [44] and OXE [45]. Additionally, we want to emphasize ecologically valid language instructions that perturbations may deviate from, which could involve new scene perturbations (e.g., new object instances), language perturbations (e.g., changing cultural contexts), or different cost types (e.g., labor costs). We assumed the existence of a test set that is representative of the test manifold, and that contrast set perturbations will generate instances within the domain. Unlike other works, we do not aim to estimate performance outside this domain. Typically, test sets are not well-defined in advance and future work can explore which perturbations can allow experimenters to efficiently evaluate more open-ended robot systems. Now that our work has shown the utility of perturbation strategies for efficiently evaluating robot policies, future work can consider automatic approaches to selecting which perturbations to apply and for generating new test instances for experimenters, potentially using large language models.
@@ Acknowledgments
This work was supported in part by a grant from the Army Research Lab (ARL) Army AI Innovations Institute (A2I2), award number W911NF-23-2-0010. Rohan Gupta was supported by a USC Provost’s Undergrad Research Fellowship during a portion of this work. The authors would like to thank Elle Szabo for building the early infrastructure for the physical robot experiments.
Appendix A Contrast Sets for Robotics
Overall, our work investigates whether contrast sets are able to estimate the sample mean performance of a standard evaluation set with a similar or lower cost. We do not make estimates of the true population mean of the performance (i.e. run the experiment a large number of times so that the performance converges to a single true mean). Instead, we focus on sample mean performance comparisons because experimenters can realistically only estimate the sample mean and not the true population mean.
Algorithm 1 Contrast set with perturbations
1:Input: Test set $X$, budget $K$, perturbation functions $\mathcal{P}$
2:$c_{total}\leftarrow 0$
3:for $x\in X$ do
4:     for $\delta\in\mathcal{P}$ do
5:         $c\leftarrow\text{cost}(x^{\prime},x)$
6:         $x^{\prime},\hat{y}\leftarrow f(\delta(x))$triangleright $x^{\prime}$ is the resultant state
7:         $c_{total}\leftarrow c_{total}+c$
8:         if $c_{total}>K$ then
9:              End
10:         end if
11:     end for
12:end for
Appendix B Language-Table Simulator Experiments
@@ B.3 Additional Results
Appendix C VLN-CE Evaluation on a Physical Robot